OpenStudy (anonymous):

Differential Equations! Currently I have: ln y = 3 ln(x^2+4) and it is given that y = 32 when x = 0. How do I go on from here, I'm not too sure! Fan and Medal

OpenStudy (gorv):

ln y= ln((x^2+4)^3)

Miracrown (miracrown):

Moreover, in order to apply this initial condition y = 32 when x = 0 we need an unknown constant I don't see one in the equation Can you give me the initial problem statement? So I can dot all the i's and cross all the t's so to speak.

OpenStudy (gorv):

property of log

Miracrown (miracrown):

No room for an initial condition there, know what I mean?

OpenStudy (anonymous):

I'm not so sure what you mean Miracrown...

OpenStudy (gorv):

y=(x^2+4)^3

OpenStudy (gorv):

thats what we got here

OpenStudy (anonymous):

I have ln y = 3 ln(x^2+4) + ln(32/64)

Miracrown (miracrown):

Hm, give me the initial DE and I'll show you

OpenStudy (gorv):

that is you got after solving

OpenStudy (gorv):

initially what was the Q

OpenStudy (anonymous):

Sure, but I never learnt anything about dotting stuff. It is question 3 here http://papers.xtremepapers.com/CIE/Cambridge%20International%20A%20and%20AS%20Level/Mathematics%20(9709)/9709_w12_qp_33.pdf

OpenStudy (anonymous):

Question 4 I mean

Miracrown (miracrown):

that's a logarithmic equation

Miracrown (miracrown):

Ah, okay, looks like you made good progress

OpenStudy (gorv):

which Q??

OpenStudy (anonymous):

After taking the logs out I get y = 1/2(x^2+4)^3 but the answers say y = 1/2 (x^2+4) So apparently there's no cubed...

OpenStudy (anonymous):

Quetsion 4

OpenStudy (gorv):

wait

Miracrown (miracrown):

but the right hand side doesn't need a logarithmic anti-derivative

OpenStudy (gorv):

\[\frac{ dy }{ y }=\frac{ 6x*dx }{ x^2+4 }\]

OpenStudy (gorv):

now we need to integrate it

OpenStudy (anonymous):

Yeah got that which is: ln y = 3 ln(x^2+4)

OpenStudy (gorv):

ln y =3*ln(x^2+4) + c

OpenStudy (gorv):

no no we also add a integration constant

OpenStudy (gorv):

c

OpenStudy (anonymous):

Oh yeah forgot about that :/ Then we put in the values for y = 32 and x = 0 right?

OpenStudy (gorv):

ln y =3*ln(x^2+4) + c now in this put y=32 and x=0

OpenStudy (gorv):

and calculate c

OpenStudy (anonymous):

c = ln (32/64)

OpenStudy (gorv):

ln 32=3*ln(0+4)+c ln32=3*ln4 +c ln32=ln4^3 + c ln32=ln64 +c ln 32-ln 64=c c=ln(32/64)=ln(1/2)=ln 1 - ln 2 =0- ln 2

OpenStudy (gorv):

yeahhhhh

OpenStudy (anonymous):

So what do we do next? We've got ln y = 3 ln (x^2+4) + ln(32/64)

OpenStudy (gorv):

yepppppppppppppppppp

OpenStudy (gorv):

lny=ln(x^2+4)^3+ln (1/2) 32/64=1/2 and ln(1/2)=ln 1 - ln 2 =-ln 2 as ln 1 =0 ln y =(ln (x^2+4)^3-ln 2 ln y = ln[{(x^2+4)^3}/2] now ln on oth side will get cancelled

OpenStudy (anonymous):

so the final answer is y = 1/2 (x^2+4)^3?